Operations Report – January 5th

Crew 186 Operations Report 05JAN2018

SOL:05

Name of person filing report: M. Grande

Non-nominal systems: Generator power system

Notes on non-nominal systems: Generator has failed a couple times at start-up (today and yesterday) and reads “under speed” or “under voltage”. I think this might be the cold, as I continue to charge the batteries all day (or all night). The generator will run on the new 9/17 battery tonight, which was charging all day, but it took 4 attempts to start it.

Generator

(hours run): 15.8h

Battery

switched from old (10/17) to new (9/17) at 8:45am

Generator

turned off, charging battery at 12:30pm

Generator

turned on at 6:00pm

Solar—

SOC

@

8:45am : 68%

@

12:30pm : 100%

@

6:00pm : 65%

Diesel:

55%

Propane:

34%

Ethanol

Free Gasoline (5 Gallon containers for ATV): 8.65 Gallons

Water

(trailer): 10 Gallons

Water

(static): 318 Gallons

Trailer

to Static Pump used: No

Water

(loft) – Static to Loft Pump used: Yes

Water

Meter: 129097.7 Gallons

Toilet

tank emptied: No

ATVs

Used: 350.2

Oil

Added? No

ATV

Fuel Used: 0.1 Gallons

#

Hours the ATVs were used today: 00:40 hours

Notes

on ATVs: ATVs were nominal.

Deimos

rover used: No

Hours:

97.4

Beginning

charge: 100%

Ending

charge:

Currently

charging: Yes

Sojourner

rover used: Assigned to director only.

Hours:

Director discretional hours

Beginning

charge:

Ending

charge:

Currently

charging: Maybe

Spirit

rover used: Yes

Hours:

9.3

Beginning

charge: 100%

Ending

charge: 87%

Currently

charging: No

Opportunity

rover used: Yes

Hours:

4.9

Beginning

charge: 100%

Ending

charge: 66%

Currently

charging: Yes

Curiosity

rover used: No

Hours:

3.1

Beginning

charge: 100%

Ending

charge:

Currently

charging: No

HabCar

used and why, where? No

General notes and comments: No EVA for me today, but I served as HabCom running the airlock depressurization/ repressurization cycles, writing down times, and sending Morse Code transmissions to our NAV, Justin, through his homemade Yagi antenna. It was my intent to hop on the stationary bike left by Crew 184 and get some exercise, but instead I spent most of the day writing up some reports and scripts for my virtual reality (VR) experiment… and enjoying some fresh, warm herb bread made in the bread maker by GEO Cesare! Yum!

Summary of internet: All nominal

Summary of suits and radios: All nominal

Summary of Hab operations: All nominal

Summary of GreenHab operations: All nominal

Summary of ScienceDome operations: All nominal

Summary of RAM operations: Not Operational

Summary of health and safety issues: Crew is Healthy

Questions, concerns and requests to Mission Support: None

Regards,

Melanie Grande, Crew 186 Engineer

Journalist Report – January 5th

[Sol 5]

Temperatures have plummeted and the night left a thin dusting of frost that lingered in the shadows throughout the day. Last night the team planned one of our most complex EVAs yet. One that would require the maximum compliment of 5 crewmembers on EVA at once. The goal: locate elusive hematite “blueberries” (named for their rounded shape and color) that could indicate the presence of liquid surface water in the distant Martian past.

The convoy travelled south to revisit areas explored on EVA #2. This time the team focused our search around a possible ancient stream bed. But alas, the search was unsuccessful. Still, the EVA traversed a variety of terrain and allowed all members of the EVA to gain experience identifying the different geologic formations. Back at the hab, our health and safety officer, Sam Albert, has begun cataloguing microbial samples from around the hab to better understand the microbial environment.

I am also happy to report that last night’s pizza was a hit, if not only due to Cesare’s magic touch that saved the dough. The crew was able to fit in some astronomical observations of the double Martian moonrise before relaxing over a television show. Tomorrow we will continue these leisure activities and have dedicated the day to relaxation to mark the halfway point of our mission.

Justin Mansell, MDRS Crew 186 Journalist

P.S. Photos attached. Photo of the day: 05Jan2018 Building a microbe catalogue.jpg

Science Report – January 5th

Science: Geology

EVA #5 revisited the location to the East of Greenstone Rd, in search of the elusive hematite-coated “blueberries”. Unfortunately, though spectra showed traces of hematite mixed with clay, no spherules were located. Nonetheless, the long EVA was satisfactory since it went through a variety of terrains, including stream beds, dunes, and various layers of sandstone and conglomerate formations. On the way back, the crew stopped at the Kissing Camel Ridge, where -among the sandstone and mudstone layers of the Morrison formation- the crew found boulders collapsed from the Dakota sandstone and conglomerates that top the Morrison formation.

Cesare Guariniello, PhD

 

EVA Report – January 5th

Author: Max Fagin

Purpose of EVA: Geological sampling of the Greenstone Rd. area and continuing the search for hematite blueberries

Destination: South of Greenstone Rd

UDM27 Coordinates:

520000 E: 424850 N

Participants: Max Fagin (Commander), Kshitij Mall (Executive Officer), Cesare Guariniello (Geologist), Justin Mansell (Journalist), Mark Gee (GreenHab Officer).

Narrative: We reformatted the EVA after suiting up and a radio conversation with Shannon to be an exclusively geological EVA, and delay the navigation component to a later EVA (we previously misunderstood the EVA rules about crew members separating into teams after departing the hab, and how this would technically constitute two separate EVAs). We departed the Hab at 11:10 and arrived on site at 11:48. We walked a long counterclockwise loop around the area searching the ground and occasionally stopping for spectral samples with Cesar’s portable spectrometer, and for the first time, with the portable laser spectrometer that had been generously lent to us by NASA Ames. Despite spending 3 hours on site and seeing many interesting geologic features, we did not find any of the hematite blueberry deposits that are reported to be in this area. We stopped at the Kissing Camels feature on the way back to sample a layer of Dakota Sandstone from the Cretaceous period and returned to the Hab at 14:30.

I expected a 5 person EVA to only be ~20% harder than our more typical 4 person EVA, but in reality, my impression is that they are ~50% harder. In addition to requiring multiple press/depress cycles, the 5th crew member requires an extra vehicle and more room than is easily available in the suit up area. We will probably stick to the conventional 4 crew EVAs in the future. That really does seem to be the perfect number.

Sol Summary Report – January 5th

Crew 186 Sol 5 Summary Report 05JAN2018

Sol 5

Summary Title: The Hunt for Blueberries and Microbial DNA

Author’s name: Max Fagin

Mission Status: Excited for our first scheduled R&R

Sol Activity Summary: We awoke today to Hans Zimmer’s “Elysium” from the movie Gladiator. Clouds had rolled in as forecasted, preventing the solar panels from being used for daytime power. The generator had failed during the night, and while it was easy to restart, it took a few attempts before it kept running for more than ~10 seconds. We suspect it simply had to warm up, but our reducing the hab power load before turning on the generator may have also helped. After breakfast, we prepared for our first 5 crew EVA to Greenstone Rd to continue the hunt for the hematite blueberries that had eluded us on our first and second EVA. After suit up and a brief radio discussion with Shannon, we decided to drop the navigation component of the EVA and focus on pure geology. Unfortunately, although we spent almost 3 hours on site with 5 pairs of eyes, and obtained multiple geological samples and spectra from around the area (including from the laser spectrometer generously lent to us by NASA Ames) we were not able to find any Hematite Blueberry formations. This evening, we got our first positive results from Sam’s microbe DNA experiment after swabbing around the hab. The PCR machine will continue to run throughout the night, and we eagerly await tomorrow’s results!

Look Ahead Plan: To commemorate reaching the midpoint of our mission, tomorrow will be a light duty day of Hab upkeep and catching up on photo/video uploads and emails, followed by some R&R. No EVAs are planned for tomorrow.

Anomalies in work: Generator failures

Weather: -2C, Winds E at 5 MPH, visibility 10 miles,

Crew Physical Status: Healthy

EVA: Greenstone Rd, 3.0 hrs, Geological sampling and search for hematite blueberries

Reports to be filed: Sol Summary, Journalist Report, Operations Report, Science Report (Geology), Greenhab Report, EVA Report

Support Requested: The generator failures are concerning, as we have no way to detect them unless someone is outside or in the science dome to hear the generator turn off. We can schedule a round of hourly checks, but it would interfere with our sleep schedule. Any recommendations on ways to prevent/detect the generator failures?

Our water conservation efforts are probably not going to be sufficient, as we are projecting a ~1 sol deficit in our initial allotment of 550 gallons. Of this, 40% is going to Hab usage for drinking, washing and hygiene, 7% is watering our crop experiments, and 53% to watering the crops that were present in GreenHab at handoff. Would it be possible to schedule a water resupply so that we can ensure no interruption to the crops?

Tomorrow we plan to catch up on the photo/video backlog that we have developed. We have heard conflicting reports on upload limits during regular hours. Is it truly unlimited, or is that a mistake?

Greenhab Report – January 5th

GreenHab Report

Mark Gee

05Jan2018

Environmental control:

Heating

Shade Cloth On

Working Hour: 06:30 PM
Inside temp at working hour: 16 C
Outside temp during working hours: -4 C
Inside temperature high: 29 C
Inside temperature low: 14 C
Inside humidity: 30 %RH

Inside humidity high: 32 %RH
Inside humidity low: 16 %RH

Hours of supplemental light:

For the crops 05:00 to 11:59 PM

Changes to crops: Transplanted tomatos are growing well. Kale sprouts have grown their first leaves. Cucumbers are flowering profusely and baby cucs abound. Watered all crops with MiracleGrow at a concentration of 1tsp/gal as recommended by the supplier. Next fertilization date is 19Jan2018.

Daily water usage for crops: 8 gallons

Time(s) of watering for crops: 06:30 PM

Research observations: Moringa experiment is growing well. Microgreens have radicles and cotyledons that are turning green.

Changes to research plants: I accidentally watered two of the tomato plants with the MiracleGrow, so I moved these plants into normal crop production.

Aquaponics: Not Functional

Narrative: Clouds obscured the sun all day and kept the Green Hab at a moderate temperature. I was on an EVA all day and did not have much time to spend inside. Tomorrow I will be seeding several vegetables for future crews to enjoy.

Support/supplies needed: None